### v1

@@ 3.1 General Weighted Obstacles in Oracle Model
Compute the intersection graph in the homology cover $\overline{G}$ of $\mathcal{C}$. For each $\gamma\in\mathcal{C}$, compute the shortest-path from $(\gamma,-)$ to $(\gamma,+)$ in $\overline{G}$. The smallest such path induces the solution by LABEL:lemma:sp. Each shortest-path can be computed in $O(k+|\mathcal{C}|\log|\mathcal{C}|)$ time by Dijkstra’s algorithm with a Fibonacci heap [clrs], and there are $O(|\mathcal{C}|)$ total such paths.
Note that theorem 3.1 matches the result of [CabelloG16], although the algorithm is much simpler. When $k\in\Theta(|\mathcal{C}|^{2})$ (i.e., dense intersection graphs), this algorithm runs in $O(|\mathcal{C}|^{3})$ time, which could also be obtained by running Floyd-Warshall for APSP instead of Dijkstra. It is not known that the general form of weighted APSP can be solved in “truly” subcubic time (see LABEL:subsec:apsp). However, since the weights are applied to vertices, not edges, we can obtain a faster algorithm via the results from the discussion in LABEL:subsec:apsp:
@@ 3.2 General Unweighted Obstacles in Oracle Model
We are also interested in the $(s,t)$ point-separation problem for unit weights, which corresponds to shortest-paths for unit weights. In this case, we can obtain faster algorithms:
@@
For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$ that have $k$ pairwise intersections, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ time, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
@@ 3.3 General Weighted Obstacles in Arrangement Model
  • Compute the SSSP tree from each obstacle in $\overline{H}$.
@@
Let $k=\min(|\mathcal{C}|,|V(D)|)$. Each SSSP computation can be performed in $O(m+k\log k)$ time, and $k$ such computations suffices to solve the problem.
@@ 3.4 Restricted Obstacle Classes without Weights
[…]
@@
Run APSP for the intersection graph in the homology cover via LABEL:thm:unweighted-intersection-apsp.
To use this result, we need to algorithms for static intersection in the homology cover:
@@
[…]

### v2

@@ 3.1 General Weighted Obstacles in Oracle Model
Compute the intersection graph in the homology cover $\overline{G}$ of $\mathcal{C}$. For each $\gamma\in\mathcal{C}$, compute the shortest path from $(\gamma,-)$ to $(\gamma,+)$ in $\overline{G}$. The smallest such path induces the solution by LABEL:lemma:sp. Each shortest path can be computed in $O(k+|\mathcal{C}|\log|\mathcal{C}|)$ time by Dijkstra’s algorithm with a Fibonacci heap [clrs], and there are $O(|\mathcal{C}|)$ total such paths.
Note that Theorem 3.1 matches the result of [CabelloG16], although the algorithm is much simpler. When $k\in\Theta(|\mathcal{C}|^{2})$ (i.e., dense intersection graphs), this algorithm runs in $O(|\mathcal{C}|^{3})$ time, which could also be obtained by running Floyd-Warshall for APSP instead of Dijkstra. It is not known if the general form of weighted APSP can be solved in “truly” subcubic time (see LABEL:subsec:apsp). However, since the weights are applied to vertices, not edges, there is a known faster algorithm (also discussed in LABEL:subsec:apsp):
@@ 3.2 General Unweighted Obstacles in Oracle Model
We are also interested in the $(s,t)$ point-separation problem for unit weights, which corresponds to shortest paths for unit weights. In this case, we can obtain faster algorithms:
@@
For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ time, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
@@ 3.3 General Weighted Obstacles in Arrangement Model
  • Compute the single-source shortest-path (SSSP) tree from each obstacle in $\overline{H}$.
@@
Let $k=\min(|\mathcal{C}|,|V(D)|)$. Each SSSP computation can be performed in $O(m+k\log k)$ time, and $k$ such computations suffice to solve the problem.
@@ 3.4 Restricted Obstacle Classes without Weights
For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$. Let $\overline{\mathcal{C}}$ be the set of $2|\mathcal{C}|$ “mapped” obstacles in the homology cover. Let $\text{SI}(n,m)$ (“static intersection”) be the time complexity for checking if each of $n$ different objects in $\overline{\mathcal{C}}$ intersects any object in some subset $C\subset\overline{\mathcal{C}}$ of size $m=|C|$. Assume $\text{SI}(n,m)$ is super-additive, so that $\text{SI}(n_{1},m_{1})+\text{SI}(n_{2},m_{2})\leq\text{SI}(n_{1}+n_{2},m_{1}+m_{2})$. Then $(s,t)$ point-separation over $\mathcal{C}$ can be solved in $O(n\text{SI}(n,n))$ time.
@@
Run APSP for the intersection graph in the homology cover via the algorithm of Chan [ChanS19, Theorem 2] (see further discussion in LABEL:subsec:apsp).
To use this result, we need efficient algorithms for static intersection in the homology cover:
@@
For points $s$ and $t$, the following values of $\text{SI}(n,n)$ hold for restricted obstacle types in the homology cover:
Obstacle Class $\text{SI}(n,n)$